OpenStudy (anonymous):

Jodie uses her college printer to print pages at the rate of $0.06 per page. She decides to rent a printer for $80 a year. The cost of printing using the rented printer is $0.02 per page. Here is my inequality: 80 + 0.02p < 0.06p How many pages should Jodie print in a year to justify renting the printer? Show your work.

OpenStudy (anonymous):

I need help solving it & I need help with the problem under the inequality.

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

Well, it says justify.. So I would assume less or equal amount. I am not really sure

OpenStudy (phi):

oh, never mind.. I was confused. Your inequality is fine. Now we need to solve it.

OpenStudy (phi):

I would add -0.02p to both sides as a first step.

OpenStudy (anonymous):

Okay & Yes I need help with that.. I think you have to subtract 0.02p but I have no idea .-.

OpenStudy (anonymous):

Okay.

OpenStudy (phi):

what do you get ?

OpenStudy (anonymous):

0.04p??

OpenStudy (phi):

what do you get for the whole equation ?

OpenStudy (anonymous):

80<0.04p

OpenStudy (phi):

yes. 80 + 0.02p < 0.06p 80 +0.02p - 0.02p < 0.06p - 0.02p 80 < 0.04p

OpenStudy (phi):

what should you divide both sides by to get "p" by itself?

OpenStudy (anonymous):

0.04..

OpenStudy (phi):

yes

OpenStudy (anonymous):

OKay & I got 2000

OpenStudy (anonymous):

so p=2000

OpenStudy (phi):

you should write out the whole thing: 2000< p which "says" p is bigger than 2000 and that means if you print more than 2000 pages, the rental is cheaper.

OpenStudy (anonymous):

OKay can you help me with one more??

OpenStudy (phi):

?

OpenStudy (anonymous):

Part A: Solve C = 5 over 9 (F - 32) for F. (4 points) Part B: Determine the value of F when C = 8 °C. (2 points) Part C: Solve –np – 80 > 60 for n. Show your work. (4 points)

OpenStudy (phi):

any ideas ?

OpenStudy (anonymous):

Not really.. I don't remember how to do this .-.

OpenStudy (phi):

you could, but I would multiply both sides by the reciprocal of 5/9

OpenStudy (anonymous):

OKay so it would like this C = 5/9 (F-32) (5/9)???

OpenStudy (phi):

yes, but with equations you do the same thing to both sides

OpenStudy (anonymous):

5/9 (5/9) (F-32) (5/9)

OpenStudy (phi):

where is the C ? and the = sign ?

OpenStudy (anonymous):

Sorry I forgot it.. C= 5/9 (5/9) (F-32) (5/9)

OpenStudy (anonymous):

Would I multiply it with C?

OpenStudy (phi):

ok, pay attention: \[ C = \frac{5}{9} (F-32) \] the reciprocal of a fraction is the "flipped fraction"

OpenStudy (phi):

write 9/5* on each side

OpenStudy (phi):

\[ \frac{9}{5}C = \frac{9}{5}\cdot \frac{5}{9} (F-32) \]

OpenStudy (anonymous):

9/5 * C= 9/5 * 5/9 (f-32)

OpenStudy (phi):

and what is 9/5 * 5/9 ?

OpenStudy (anonymous):

One second

OpenStudy (anonymous):

no no 1

OpenStudy (phi):

yes, the idea is that the 9's cancel and the 5's cancel. (That's the reason I picked it) what do we have now ?

OpenStudy (anonymous):

Just this : C = 1(F-32)

OpenStudy (anonymous):

Can we distribute now??

OpenStudy (phi):

wait. we started with \[ \frac{9}{5}C = \frac{9}{5}\cdot \frac{5}{9} (F-32) \]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

oh so its 9/5 C = (f-32)

OpenStudy (phi):

yes. you can remove the parens (they are not doing anything) 9/5 C = f -32 what should you add to both sides?

OpenStudy (phi):

you should get \[ F= \frac{9}{5} C +32 \]